On January 1, 2015, a machine was purchased for $101,700. The machine has an estimated salvage value of $6,780 and an estimated useful life of 5 years. The machine can operate for 113,000 hours before it needs to be replaced. The company closed its books on December 31 and operates the machine as follows: 2015, 22,600 hrs; 2016, 28,250 hrs; 2017, 16,950 hrs; 2018, 33,900 hrs; and 2019, 11,300 hrs.Please show calculations for all problems below.

1) Compute the annual depreciation charges over the machine's life assuming a December 31 year-end for each of the following depreciation methods:

a) Straight-Line Method $___

b) Activity Method

2015 $__ 2016 $__ 2017 $__ 2018 $__ 2019 $__

c) Sum-of-the-Years' Digits Method

2015 $__ 2016 $__ 2017 $__ 2018 $__ 2019 $__

d) Double-Declining-Balance Method

2015 $__ 2016 $__ 2017 $__ 2018 $__ 2019 $__

2) Assume a fiscal year-end of September 30. Compute the annual depreciation charges over the asset’s life applying each of the following method:

Year Straight-Line Method Sum-of-the-years-digits Method Double-declining Method

2015 $__ $__ $__

2016 $__ $__ $__

2017 $__ $__ $__

2018 $__ $__ $__

2019 $__ $__ $__

2020 $__ $__ $__
Answer #1

Answer 1:

Table 1:

Depreciation SLM Activity method Sum of years' digit method Double declining balance method
2015 $     18,984.00 $       18,984.00 $                31,640.00 $             40,680.00
2016 $     18,984.00 $       23,730.00 $                25,312.00 $             24,408.00
2017 $     18,984.00 $       14,238.00 $                18,984.00 $             14,644.80
2018 $     18,984.00 $       28,476.00 $                12,656.00 $               8,786.88
2019 $     18,984.00 $          9,492.00 $                  6,328.00 $               5,272.13
Total $ 94,920.00 $ 94,920.00 $ 94,920.00 $ 93,791.81

Answer 2:

Table 2:

Depreciation SLM Sum of years' digit method Double declining balance method
30-Sep-15 $        9,492.00 $       15,820.00 $                20,340.00
30-Sep-16 $     18,984.00 $       28,476.00 $                32,544.00
30-Sep-17 $     18,984.00 $       22,148.00 $                19,526.40
30-Sep-18 $     18,984.00 $       15,820.00 $                11,715.84
30-Sep-19 $     18,984.00 $          9,492.00 $                  7,029.50
30-Sep-20 $        9,492.00 $          3,164.00 $                  2,636.06
Total $ 94,920.00 $ 94,920.00 $ 93,791.80

Working note (Table 3): Values for Table 2

Depreciation SLM Sum of years' digit method Double declining balance method
01-Jan-15 to 30-Sep-15 $        9,492.00 $       15,820.00 $                20,340.00
01-Oct-15 to 31-Dec-15 $        9,492.00 $       15,820.00 $                20,340.00
01-Jan-16 to 30-Sep-16 $        9,492.00 $       12,656.00 $                12,204.00
01-Oct-16 to 31-Dec-16 $        9,492.00 $       12,656.00 $                12,204.00
01-Jan-17 to 30-Sep-17 $        9,492.00 $          9,492.00 $                  7,322.40
01-Oct-17 to 31-Dec-17 $        9,492.00 $          9,492.00 $                  7,322.40
01-Jan-18 to 30-Sep-18 $        9,492.00 $          6,328.00 $                  4,393.44
01-Oct-18 to 31-Dec-18 $        9,492.00 $          6,328.00 $                  4,393.44
01-Jan-19 to 30-Sep-19 $        9,492.00 $          3,164.00 $                  2,636.06
01-Oct-19 to 31-Dec-19 $        9,492.00 $          3,164.00 $                  2,636.06
Total $     94,920.00 $       94,920.00 $                93,791.81

Formulae for Table 1:

Depreciation SLM Activity method Sum of years' digit method Double declining balance method
2015 =(101700-6780)/5 =(101700-6780)/113000*22600 =(101700-6780)/15*5 =101700*40%
2016 =(101700-6780)/5 =(101700-6780)/113000*28250 =(101700-6780)/15*4 =101700*60%*40%
2017 =(101700-6780)/5 =(101700-6780)/113000*16950 =(101700-6780)/15*3 =101700*60%*60%*40%
2018 =(101700-6780)/5 =(101700-6780)/113000*33900 =(101700-6780)/15*2 =101700*60%*60%*60%*40%
2019 =(101700-6780)/5 =(101700-6780)/113000*11300 =(101700-6780)/15*1 =101700*60%*60%*60%*60%*40%

Formulae for Table 2:

Depreciation SLM Sum of years' digit method Double declining balance method
30-09-2015 9492 15820 20340
30-09-2016 =9492+9492 =15820+12656 =20340+12204
30-09-2017 =9492+9492 =12656+9492 =12204+7322.4
30-09-2018 =9492+9492 =9492+6328 =7322.4+4393.44
30-09-2019 =9492+9492 =6328+3164 =4393.44+2636.06
30-09-2020 9492 3164 =2636.06

Formulae for Table 3:

Depreciation SLM Sum of years' digit method Double declining balance method
01-Jan-15 to 30-Sep-15 =((101700-6780)/5)/2 =((101700-6780)/15*5)/2 =(101700*40%)/2
01-Oct-15 to 31-Dec-15 =((101700-6780)/5)/2 =((101700-6780)/15*5)/2 =(101700*40%)/2
01-Jan-16 to 30-Sep-16 =((101700-6780)/5)/2 =((101700-6780)/15*4)/2 =(101700*60%*40%)/2
01-Oct-16 to 31-Dec-16 =((101700-6780)/5)/2 =((101700-6780)/15*4)/2 =(101700*60%*40%)/2
01-Jan-17 to 30-Sep-17 =((101700-6780)/5)/2 =((101700-6780)/15*3)/2 =(101700*60%*60%*40%)/2
01-Oct-17 to 31-Dec-17 =((101700-6780)/5)/2 =((101700-6780)/15*3)/2 =(101700*60%*60%*40%)/2
01-Jan-18 to 30-Sep-18 =((101700-6780)/5)/2 =((101700-6780)/15*2)/2 =(101700*60%*60%*60%*40%)/2
01-Oct-18 to 31-Dec-18 =((101700-6780)/5)/2 =((101700-6780)/15*2)/2 =(101700*60%*60%*60%*40%)/2
01-Jan-19 to 30-Sep-19 =((101700-6780)/5)/2 =((101700-6780)/15*1)/2 =(101700*60%*60%*60%*60%*40%)/2
01-Oct-19 to 31-Dec-19 =((101700-6780)/5)/2 =((101700-6780)/15*1)/2 =(101700*60%*60%*60%*60%*40%)/2

In case of any doubt, please comment.